# Parametric Instabilities of Correlated Quantum Matter

### v1

@@ III.3 Fidelity susceptibility
Clearly, the off-diagonal component $\delta{\cal Z}_{\bf Q}$ in the perturbed Hamiltonian is finite iff $V_{\bf Q}$ is non-zero. In the limit of small perturbations $\lambda\to 0$, the parametric susceptibility and the residual transformation are proportional to one another,
$\frac{\partial\|\delta{\cal Z}_{{\bf Q}}\|}{\partial\lambda}|_{\lambda\to 0}\propto\frac{\partial\|V_{{\bf Q}}\|}{\partial\lambda}|_{\lambda\to 0},$ (27)
@@
We may now consider the fidelity of the unperturbed fluctuation vacuum $|{\bf{0}}_{\bf Q}\rangle$ with the perturbed vacuum $|{\bf\tilde{0}}_{\bf Q}\rangle$ (Appendix C),
$\left|\langle{\bf\tilde{0}}_{{\bf Q}}|{\bf 0}_{{\bf Q}}\rangle\right|=\left[{\det\left(\mathbb{1}_{N}+V_{\bf Q}V_{\bf Q}^{\dagger}\right)}\right]^{-1/4}.$ (28)
The momentum resolved fidelity susceptibility is then,
@@
Taken together with Eq. (27), we find that the magnitude of the parametric driving is proportional to the square root of the Fidelity susceptibility of the fluctuation vacuum. In the next section, we make this connection even more concrete with a generic example.
III.4 Single mode example
Let us demonstrate these principles on a parametric resonance of a single mode labeled $\bar{n}$ for which,
@@ III.4 Single mode example
For simplicity, we assume the mode is decoupled from all other modes, $\left[\delta{\cal Z}\right]_{n\bar{n}}\approx 0,\,\,\,\forall n\neq\bar{n}$, and we denote
@@
The formalism discussed above is general for arbitrary connectivity of the modes, yet for the salient features we explore here (and in the illustrative examples below) it suffices to consider this simpler case.
@@
consider the mode resolves susceptibilities,
$\tilde{\chi}_{F}^{\delta H}=-\frac{\partial^{2}\left|\left\langle{\bf 0}_{n,{\bf Q}}\left(H\right)|{\bf 0}_{n,{\bf Q}}\left(H+\lambda\delta H\right)\right\rangle\right|}{\partial\lambda^{2}}|_{\lambda\to 0},$ (32)
@@
$\tilde{\chi}_{F}^{\delta H}=\frac{1}{4{\cal E}_{\bar{n},{\bf Q}}^{2}}\left(\frac{\partial\left|\delta z\right|}{\partial\lambda}\right)^{2},\,\,\,\,\,\,\tilde{\kappa}^{\delta H}=\frac{1}{{\cal E}_{\bar{n},{\bf Q}}}\left(\frac{\partial\left|\delta z\right|}{\partial\lambda}\right)^{2}.$ (34)
It is no surprise that these results align perfectly with Eqs. (27) and (29). We have thus demonstrated that parametric instabilities of the correlated system directly probe the quantum fluctuations in the reference ordered state. Introduction of a parametrically resonant perturbation is a sensitive measure of the response of quantum fluctuations to this perturbation.
@@
Straightforward calculation (Appendix D) then reveals
$\delta z=-\frac{\omega+\left|\upsilon\right|}{2}\left[\frac{\left(1+r^{2}\right)^{2}}{2r}d\frac{\left|\upsilon\right|}{\omega}+i\left(1-r^{2}\right)d\phi\right].$ (38)
@@
Notably, the first term inside the parenthesis in Eq. (38) diverges in the strong squeezing limit, $r\to 0$. Thus, we uncover a dramatic effect – the more squeezed the equilibrium fluctuation vacuum is, the more susceptible it is to parametric instabilities.
@@
The exponential growth signals the instability of the ordered phase with respect to the parametric resonance of this mode. Dissipation however, may alter the eventual fate of the system. Consider linear dissipation, independent of “photon” number, inducing a finite collective-mode lifetime $\tau$. A necessary condition for the parametric instability is $\left|\delta z\right|>\tau^{-1}$. We note that one source of dissipation expected for the systems we discuss, namely interactions with the particle-hole excitation continuum (e.g., the so-called Landau damping), is absent for the specific scenarios considered above.
@@ IV Case studies
In the following we will explore several possible realizations of parametrically driven correlated materials, each highlighting different facets of our general discussion above. Practically speaking, one is interested in the precise strength of the perturbation-induced parametric term. Given the expected exponential growth induced by it, it is reasonable to compare the energy of the two-boson term $\delta z$ to the energy of the driven collective mode ${\cal E}_{\bar{n},{\bf Q}}$. This is an effective measure of the number of evolution cycles of that mode before the parametric amplification is evident.
We define the parametric efficiency $\eta_{{\rm par}.}^{\left(p\right)}$ with respect to a small change $\delta p$ in the parameter $p$ [in direct analogy to Eq. (2), where the efficiency is unity],
With regards to our general single-mode result Eq. (38), The parametric efficiency corresponding to the $\upsilon$ and $\omega$ variables is expressed entirely in terms of the squeezing parameter,
which diverges as the squeezing becomes more extreme, $r\to 0$. In the examples below, we will generally find that the parametric efficiency is of order unity, unless the modulated parameter drives a phase transition of the microscopic Hamiltonian. In that case the efficiency may diverge, reflecting an expected divergence of the corresponding fidelity susceptibility.
We note that linear dissipation would compete with the driving, leading to reduced effective efficiency $\tilde{\eta}=\eta\left(1-\frac{{\cal E}_{\bar{n},{\bf Q}}}{Q\delta_{z}}\right)$, where $Q$ is the quality factor of the driven mode.
IV.1 Quantum Hall double layer
We first consider the well-studied example of a double layer quantum Hall system at total odd integer filling $\nu=2n+1$ [81, 82]. Both layers experience equal external chemical potential, such that in the absence of interactions and interlayer tunneling both are at filling factor $n+\frac{1}{2}$. Before introducing the specifics of this model, we note that the analysis below easily captures the salient features of many strongly-interacting (or flat-band) double layers, expected to show a similar (or generalized) excitonic ordered state, including electron-hole double layers [83], graphene or bilayer graphene double layers [84, 35], TMD double layers [85, 36, 37, 38], and moiré systems [86, 87].
@@ IV.1 Quantum Hall double layer
Our reference ordered state may thus be characterized by the matrix ${\cal O}=\ell_{x}$, such that the relevant fluctuation generators are readily found to be $\left\{{{\cal M}^{i}}\right\}=\ell_{y},\,\ell_{z}$. These identifications allow us to employ the scheme outlined in Sec. III in a straight forward manner, which we carry out in detail in Appendix E. Focusing our attention on the lowest lying Goldstone mode mentioned above, we recover the following effective single-mode parameters,
@@
At small momenta, the spectrum of this mode is ${\cal E}_{{\bf Q}}=\sqrt{\left({\cal E}_{{\bf 0}}^{{\rm QH-2\ell}}\right)^{2}+\left(\Delta_{{\rm sas}}\rho_{s}^{0}+E_{{\rm easy}}\frac{\rho_{s}^{0}-\rho_{s}^{z}}{2}\right)\left|{\bf Q}\right|^{2}}$, with the gap
@@
As expected, for $\Delta_{\rm sas}=0$, the collective mode becomes gapless and has linear dispersion.
As an example, let us consider driving the lowest-energy ${\bf Q}=0$ mode, i.e., one modifies the parameters of the system at a frequency $\omega_{d}=2{\cal E}_{\bf 0}^{\rm QH-2\ell}$. The strength of the parametric term is obtained by employing Eq. (38),
@@
where the last approximation is reasonable in most experimental systems, where the tunneling energy is smaller by orders of magnitude compared to the interaction anisotropy energy scale. The upshot of Eq. (49) is that parametric driving of the correlated double-layer system may be induced by any perturbation that modifies the energetic ratio between the interlayer tunneling and the capacitive anisotropy. One direct way to achieve this is to modulate the tunneling barrier between the layers (e.g, electrostatically or via dynamic piezoelectrical modulation), Which directly impacts the interlayer tunneling strength $\Delta_{\rm sas}$.
An alternative, more experimentally accessible approach is to modulate a gate-voltage controlled displacement electric field. Applied perpendicularly to the double-layer structure, it will induce layer polarization, i.e., a finite $\langle\ell_{z}\rangle$ component in the order parameter, rendering the fluctuations “less squeezed” in the in-plane direction, see Fig. 3a. (We note that an in-plane magnetic field may lead to a related effect, where the which-layer symmetry is spontaneously broken [89].) With a high enough displacement field, the in-plane coherence is destroyed and the system becomes layer polarized. Crudely speaking, we can emulate this change by an effective change of the anisotropy energy.
Suppose that in the vicinity of the transition this anisotropy depends on the appropriate gate-voltage $V$ as $E_{\rm easy}\propto V^{\alpha}$. It is straightforward to extract
$\eta_{{\rm par.}}^{\left(V\right)}=\alpha\eta_{{\rm par.}}^{\left(E_{{\rm easy}}\right)}.$ (51)
@@
Drawing from experimental results in relevant systems [90, 91], the transition has a width of order $\sim 10$ mV. This suggests that in the vicinity of the transition a modulation of several mV of the voltages applied are sufficient to induce a parametric instability, which becomes apparent after a few ($\lesssim 10$) driving cycles.
@@
We note that the collective mode Hamiltonian parameters depend here on the quantum geometry of the projected wavefunctions through the $f^{2}$ term in (47) (see also Appendix E and Ref. [62], relating the stiffness to quantum geometry]). Such dependence arises quite ubiquitously, which we demonstrate in the next case study. The ability to temporally modulate the quantum geometric properties of correlated electrons is thus a valuable resource in coherently driving such collective excitations.
@@
Figure 3: Driving and probing quantum Hall double layer collective excitations (Sec. IV.1). (a) Schematic Bloch sphere of the ordered phase with zero (green) and finite (purple) displacement field. Since the equator hosting the order parameter is pushed upward on the sphere when $\langle\ell_{z}\rangle$ is finite (i.e., the system develops layer polarization), the quadrature squeezing is different. Hence, parametric instability is enabled. (b) When the system is parametrically driven, an oscillatory perpendicular electric field is induced in the system. Interlayer tunneling spectroscopy is thus expected to display coherent peaks at half the driving frequency (alongside the interlayer-coherence-associated zero-bias peak). (c) In a Josephson-analogous geometry (weak links denoted by dashed black lines), the relative phase of the condensate between the driven part (driving represented by fuzzy orange oval) and the static part would result in an ac counterflow current $I\left(t\right)$ dominated by a $\omega_{d}/2$ harmonic.
@@ IV.2 Complex orders in a flat band model
The single-particle Hamiltonian of the two-dimensional system we consider here has valley $\tau_{i}$, pseudo-spin (or sub-lattice) $\sigma_{i}$, and orbital $\rho_{i}$ degrees of freedom (all Pauli matrices),
$H_{0}=-t\sum_{{\bf k}}c_{{\bf k}}^{\dagger}\left[h_{x}\left({\bf k}\right)\rho_{x}+h_{y}\left({\bf k}\right)\rho_{y}-\Delta_{\sigma}\sigma_{z}\right]c_{{\bf k}},$ (52a)
$h_{x}\left({\bf k}\right)=\frac{1+\sigma_{z}\tau_{z}}{2}\cos\alpha_{{\bf k}}+\frac{1-\sigma_{z}\tau_{z}}{2}\cos\beta_{{\bf k}},$ (52b)
$h_{y}\left({\bf k}\right)=\frac{1+\sigma_{z}\tau_{z}}{2}\sin\alpha_{{\bf k}}+\frac{1-\sigma_{z}\tau_{z}}{2}\sin\beta_{{\bf k}}.$ (52c)
The model is related to a tunable quantum metric model introduced in Refs. [106, 107]. However, presently the momentum-dependent functions $\alpha_{\bf k}$ and $\beta_{\bf k}$, which determine the quantum geometric properties of the resultant flat bands, are related to each other by a 45-degree rotation, i.e.,
$\alpha_{{\bf k}}=\zeta\left(\cos k_{x}+\cos k_{y}\right),$ (53)
The parameter $\zeta$ characterizes a quantum geometry related length scale (in units of the unit cell), referred to henceforth as the quantum metric length [108].
The Hamiltonian Eq. (52a) has a total of eight identically flat bands: for $\Delta_{\sigma}=0$ (which we assume henceforth unless mentioned otherwise) one finds four valence bands sitting at energy t-t, and four conduction bands at +t+t. Our interest will lie in one set of these isolated bands, e.g., the valence bands. Within such set, reminiscent of the decomposition of flat bands in twisted graphene, one may define a good quantum number which further classifies the bands, ${\cal C}=\sigma_{z}\tau_{z}$. (In TBG, this would correspond to the actual Chern number associated with each of these bands, yet in our model the Berry curvature vanishes everywhere – the model is topologically trivial, yet its quantum geometry is highly non-trivial for finite values of $\zeta$.)
We examine interactions between the flat band electrons, projecting to the lowest lying four bands. The simplest density-density interaction term (that will dominate the interacting physics) $H_{\rm C}$, has the same form as in Eq. (42), except its density form factors are now given by
$\Lambda_{{\bf kk^{\prime}}}={\cal F}_{{\bf kk^{\prime}}}+{\cal D}_{{\bf kk^{\prime}}}\sigma_{z}\tau_{z},$ (55)
with ${\cal F}_{{\bf kk^{\prime}}}+{\cal D}_{{\bf kk^{\prime}}}=\cos\left(\alpha_{\bf k}-\alpha_{\bf k^{\prime}}\right)$, and ${\cal F}_{{\bf kk^{\prime}}}-{\cal D}_{{\bf kk^{\prime}}}=\cos\left(\beta_{\bf k}-\beta_{\bf k^{\prime}}\right)$. If we consider the half-filled scenario, completely filling two bands out of the lowest four, it is straightforward to demonstrate that on the level of Hartree-Fock, the ground state manifold is comprised of all order parameters which commute with $\cal C$} (see Appendix F),
$\left[{\cal O}_{\rm g.s.},{\cal C}\right]=0.$ (56)
Next, let us consider two sub-leading interaction terms, whose presence favors a subset of ordered state of the manifold in Eq. (56). They will be given by
@@
with $\rho^{\hat{M}}_{\bf q}=\sum_{\bf k}c_{{\bf k+q}}^{\dagger}\Lambda_{{\bf k+q,k}}\hat{M}c_{{\bf k}}$, and $u^{\rm K/T}_{\bf q}<0$ such that each interaction term favors the order parameters in its corresponding $\hat{M}$ matrices. For the sake of simplicity, we will consider these secondary interactions to be short-ranged, i.e., momentum independent. We then employ the following definitions, $u=-\left(u^{\rm K}+u^{\rm T}\right)>0$, $u_{{\bf q}}^{{\rm T}}-u_{{\bf q}}^{{\rm K}}=-\Delta_{\rm TK}<0$ (such that positive $\Delta_{\rm TK}$ means that the so-called T-order is energetically preferable), and
$u_{\zeta}=u\left[1-J_{0}^{4}\left(2\zeta\right)\right],$ (59)
(the quantum metric length dependence originates in integration over the form factors). We will further assume that $u_{\zeta}>\Delta_{\rm TK}$. This last assumptions simply ensures that the intervalley coherent orders precipitated by these secondary interactions are energetically favorable to all others in the $\cal C$ manifold.
Before delving into the relevant fluctuations in this toy model, a brief interlude is in order, regarding the motivation for the interaction terms introduced above. In pristine twisted bilayer graphene it has been broadly established, both analytically and numerically [98, 99, 109], that the preferred order at half-filling is the so-called Kramers intervalley coherent order (KIVC), analogous to the K-order encouraged by Eq. (57). This order, however, is sensitive to the appearance of strain in realistic devices [110, 99, 111], often leading to the emergence of a rather different intervalley coherent phase, the so-called Kekulé spiral [112].
Still, even in nearly-pristine devices, with little to no strain, KIVC order is missing. Instead, a third intervalley coherent phase appears in experiments [113] – the time-reversal invariant intervalley coherent order (TIVC), analogous to the T-order preferred by Eq. (58). It has been suggested that electron-phonon interactions involving exchange of valley momentum favor the formation of the TIVC state over KIVC [103]. This scenario is emulated in our model by the choice $\Delta_{\rm TK}>0$.
Finally, alignment of the twisted graphene devices to the underlying hexagonal Boron Nitride (hBN) substrate breaks the inversion symmetry, approximately captured by a sublattice polarization term $\sim\sigma_{z}$ [114, 115], also included as a single-particle effect in our model, Eq. (52a).

### v2

@@ III.3 Fidelity susceptibility
Clearly, the off-diagonal component $\delta{\cal Z}_{\bf Q}$ in the perturbed Hamiltonian is finite iff $V_{\bf Q}$ is non-zero. In the small perturbations limit $\lambda\to 0$, the parametric susceptibility and the residual transformation are directly related to one another. For a purely off-diagonal $\delta\Lambda_{\bf Q}$, a perturbative calculation yields
$\frac{\partial\|V_{{\bf Q}}\|}{\partial\lambda}|_{\lambda\to 0}\approx\sqrt{\sum_{mn}\frac{\left|\left[\frac{\partial}{\partial\lambda}\delta{\cal Z}_{{\bf Q}}\right]_{mn}\right|^{2}}{\left({\cal E}_{m,{\bf Q}}+{\cal E}_{n,{\bf Q}}\right)^{2}}},$ (29)
@@
We may now consider the fidelity of the unperturbed fluctuation vacuum $|{\bf{0}}_{\bf Q}\rangle$ (Appendix C),
$\left|\langle{{\bf 0}_{{\bf Q}}^{\left[H\right]}}|{{\bf 0}_{{\bf Q}}^{\left[H+\lambda\delta H\right]}}\rangle\right|=\left[{\det\left(\mathbb{1}_{N}+V_{\bf Q}V_{\bf Q}^{\dagger}\right)}\right]^{-1/4}.$ (30)
The momentum resolved fidelity susceptibility,
@@
Taken together with Eq. (29), we find that the magnitude of the parametric driving is roughly proportional to the square root of the Fidelity susceptibility of the fluctuation vacuum. In the next section, we make this connection even more concrete with a generic example.
III.4 Isolated driven mode
It is instructive to consider a microscopic perturbation $\delta H$, that creates an efficient parametric excitation in a single mode, labeled by $\left(\bar{n},{\bf Q}\right)$. The periodic driving of $\delta H$ adheres to the resonant condition for this mode, i.e.,
@@ III.4 Isolated driven mode
For simplicity, we assume the mode is decoupled from all other modes, $\left[\delta{\cal Z}\right]_{n\bar{n}}\approx 0,\,\,\,\forall n\neq\bar{n}$, and we denote the induced two-boson term
@@
The formalism discussed above is general for arbitrary connectivity of the modes, yet for the salient features we explore here (and in the illustrative examples below) it suffices to consider this simpler isolated-mode case.
@@
consider the mode resolved susceptibilities,
$\tilde{\chi}_{F}^{\delta H}=-\frac{\partial^{2}\left|\left\langle{\bf 0}^{\left[H\right]}_{n,{\bf Q}}|{\bf 0}^{\left[H+\lambda\delta H\right]}_{n,{\bf Q}}\right\rangle\right|}{\partial\lambda^{2}}|_{\lambda\to 0},$ (34)
@@
$\frac{\partial\left|\delta z\right|}{\partial\lambda}=2{\cal E}_{\bar{n},{\bf Q}}\sqrt{\tilde{\chi}_{F}^{\delta H}},\,\,\,\,\,\,\frac{\partial\left|\delta z\right|}{\partial\lambda}=\sqrt{{\cal E}_{\bar{n},{\bf Q}}\tilde{\kappa}^{\delta H}}.$ (36)
Notice these results align perfectly with Eqs. (29) and (31). We have thus demonstrated that parametric instabilities of the correlated system directly probe the quantum fluctuations in the reference ordered state. Introduction of a parametrically resonant perturbation is a sensitive measure of the response of quantum fluctuations to this perturbation.
@@
Straightforward calculation(Appendix D) then reveals The dependence of the parametric (two-boson) term $\delta z$ on the infinitesimal change of fluctuation parameters $\omega$, $\upsilon$, and $\phi$,
$\delta z=-\frac{\omega+\left|\upsilon\right|}{2}\left[\frac{\left(1+r^{2}\right)^{2}}{2r}d\left(\frac{\left|\upsilon\right|}{\omega}\right)+i\left(1-r^{2}\right)d\phi\right].$ (40)
@@
It is instructive to consider the parametric efficiency defined in Eq. (8) with respect to the $\upsilon$ and $\omega$ variables. It is expressed entirely in terms of the squeezing parameter,
which diverges as the squeezing becomes more extreme, $r\to 0$. Thus, we uncover a dramatic effect – the more squeezed the equilibrium fluctuation vacuum is, the more susceptible it is to parametric instabilities.
@@
The exponential growth signals the instability of the ordered phase with respect to the parametric resonance of this mode. Dissipation however, may alter the eventual fate of the system (cf., Ref. [78]). Consider linear dissipation, independent of “photon” number, inducing a finite collective-mode lifetime $\tau$. A necessary condition for the parametric instability is $\left|\delta z\right|>\tau^{-1}$. We note that one source of dissipation expected for the systems we discuss, namely interactions with the particle-hole excitation continuum (e.g., the so-called Landau damping), is absent for the specific scenarios considered above.
@@
Figure 4: Summary of the physical examples studied in Sec. IV. (a) In Sec. IV.1 we study the parametric instability of the layer-coherent excitonic insulating phase of a correlated quantum Hall bilayer. The instability can be driven by, e.g., Floquet modulation of the interlayer tunneling amplitude (see text). The apparent squeezing between the two quadratures of the low-energy collective excitation stems from the finite energy cost associated with charging the effective two-plate capacitor. Conversely, rotating the phase of the coherent condensate is relatively much easier for the system. In the driven state, the system experiences enhanced fluctuations of interlayer charge displacement, which are similar to plasmons (yet are exactly compensated between the layers). (b) In Sec. IV.2.1 we examine a particular intervalley coherent ground state of moiré graphene. The graphic on the left panel illustrates a particular signature of this phase in an STM experiment. Focusing on one branch of the low-energy collective modes associated with this phase, we trace the squeezing in this mode to the hierarchy of competing Hartree-Fock ground states: One quadrature of the mode, corresponding to mixing a different intervalley coherent state into the order parameter, is lower in energy and thus easier for the system as compared to the conjugate quadrature (which involves sublattice polarization at the graphene scale). Upon resonant parametric driving, the inter-moiré unit-cell charge distribution, discernible by STM, fluctuates in time between sublattice polarization and sublattice-balanced patterns. (c) Finally, Sec. IV.2.2 considers sublattice-polarized order in moiré graphene. The two conjugate quadratures of the low-lying collective mode now correspond to components of two different intervalley coherent states. As a consequence, the driven state will display temporal fluctuation between a Kekulé pattern (corresponding to the T-IVC) and a trivial pattern (from the K-IVC component). Relevant drives for both cases, (b)–(c), are an electric displacement field (which modulates the quantum geometry), and non-linear phonon pumping (controls the electron-phonon coupling).
@@ IV Case studies
In the following we will explore several possible realizations of parametrically driven correlated materials, each highlighting different facets of our general discussion above. For each system, we will present the relevant correlated phase of matter, and derive the corresponding bosonic fluctuations. A summary of the phases of matter we study, the physical origin of squeezing in their collective modes, and their physical implication appears in Fig. 4. We discuss the susceptibility of these fluctuations to parametric driving using experimentally accessible probe, and show when it becomes anomaly strong by considering the appropriate parametric efficiency figure of merit [see Eq. (8)].
[…]
Twisted graphene toy model.– (Fig. 4b–c) We explore two different ordering possibilities in such systems. In a valley-coherent ground state, the parametric efficiency diverges as $\Delta_{\rm TK}^{-1}$ [Eq. (62)], where $\Delta_{\rm TK}$ is the energy difference between the ordered mean-field ground state and its closest competitor. This parameter is intimately related to electron-phonon coupling of specific modes [82], making it susceptible to periodic driving of these modes [83]. Parametrically driving the appropriate collective mode promotes a time-periodic sublattice polarization in the system (with half the driving frequency), which could be transiently detected by a second-harmonic-generation probe.
Conversely, when the electronic order corresponds to sublattice polarization, the parametric efficiency diverges in a manner inversely proportional to the distance from the phase transition to the lowest lying competing order [Eq. (65)]. We discuss how these divergences lead to detectable experimental signatures, even with small parameter modulations (e.g., $\sim$ 10 mV/nm of the applied displacement fields). For example, a time resolved STM probe would detect suppression of the ordered phase and oscillations in the real space charge distribution.
IV.1 Quantum Hall double layer (Fig. 4a)
We first consider a double layer quantum Hall system at total odd integer filling $\nu=2n+1$ [84, 85]. Both layers experience equal external chemical potential, such that in the absence of interactions and interlayer tunneling both are at filling factor $n+\frac{1}{2}$. Our analysis below easily captures the salient features of many strongly-interacting (or flat-band) double layers, expected to show a similar (or generalized) excitonic ordered state, including electron-hole double layers [86], graphene or bilayer graphene double layers [87, 35], TMD double layers [88, 36, 37, 38], and moiré systems [89, 90].
@@
Our reference ordered state may thus be characterized by the matrix ${\cal O}=\ell_{x}$, such that the relevant fluctuation generators are readily found to be $\left\{{{\cal M}^{i}}\right\}=\ell_{y},\,\ell_{z}$. The fluctuations correspond to a rotation of the coherent interlayer phase ($\ell_{y}$) and layer charge imbalance ($\ell_{z}$). These identifications allow us to employ the scheme outlined in Sec. III in a straight forward manner, which we carry out in detail in Appendix E. Focusing our attention on the lowest lying Goldstone mode mentioned above, we recover the following effective single-mode parameters,
@@
At small momenta, the spectrum of this mode is
${\cal E}_{{\bf Q}}=\sqrt{\left({\cal E}_{{\bf 0}}^{{\rm QH-2\ell}}\right)^{2}+\left(\Delta_{{\rm sas}}\rho_{s}^{0}+E_{{\rm easy}}\frac{\rho_{s}^{0}-\rho_{s}^{z}}{2}\right)\left|{\bf Q}\right|^{2}},$ (49)
with the gap at zero momentum
@@
Notice that for $\Delta_{\rm sas}=0$, the collective mode becomes gapless and has linear dispersion, since the system spontaneously (instead of explicitly) breaks the $U\left(1\right)$ symmetry.
To demonstrate and quantify the parametric instability, let us consider driving the lowest-energy ${\bf Q}=0$ mode, i.e., modifying the parameters of the system at a frequency $\omega_{d}=2{\cal E}_{\bf 0}^{\rm QH-2\ell}$. The strength of the parametric term is obtained by employing Eq. (40),
@@
where the last approximation is reasonable in most experimental systems, where the tunneling energy is smaller by orders of magnitude compared to the interaction anisotropy energy scale. The upshot of Eq. (51) is that parametric driving of the correlated double-layer system may be induced by any perturbation that modifies the energetic ratio between the interlayer tunneling and the energy associated with spontaneously creating a charge imbalance between the layers.
One direct way to achieve this is to modulate the tunneling barrier between the layers. In a double quantum well setup [92], this can be achieved by electrostatic modulation, effectively changing the wavefunction hybridization between the two wells making up the double layer. The interlayer tunneling may also be controlled using light [93, 94]. One can modulate the amplitude of the incident light on the system at the appropriate parametric resonance frequency [95], and thus create temporal oscillations of $\Delta_{\rm sas}$.
Alternatively, one may seek to modulate $E_{\rm easy}$, the easy-plane anisotropy. As can be seen from the dependence of Eq. (48) on $f\left({\bf q}\right)$, quantum geometry plays an important role. By modulating a perpendicular displacement field in, e.g., a double bilayer graphene system [96], one induces changes to the wavefunction of the active Landau levels, thus changing $E_{\rm easy}$.
@@
Drawing from the experimental results in relevant systems [97, 96], the transition out of the layer-coherent phase has a width of order $\sim 10$ mV in terms of the applied gate-voltage. This suggests that in the vicinity of the transition a modulation of several mV of the voltage applied are sufficient to induce a parametric instability, which becomes apparent after a few ($\lesssim 10$) driving cycles.
@@
We note that the dependence of the collective mode Hamiltonian parameters on the quantum geometry of the projected wavefunctions (see also Appendix E and Ref. [63], relating the stiffness to quantum geometry) arises quite ubiquitously, as we demonstrate next. The ability to temporally modulate the quantum geometric properties of correlated electrons, i.e., their Bloch wavefunctions, is thus a valuable resource in coherently driving collective excitations.
@@
Figure 5: Driving and probing quantum Hall double layer collective excitations (Sec. IV.1). (a) Schematic Bloch sphere of the interlayer-coherent ($\ell_{x}$) ordered phase. A phase-rotated coherent phase ($\ell_{y}$) and layer polarized phase ($\ell_{z}$) are schematically depicted. The squeezing of quantum fluctuations around the ordered state stems from the capacitive energy of the double layer system [Eq, (48)], manifesting an easy-plane anisotropy. (b) When the system is parametrically driven, an oscillatory perpendicular electric field is induced in the system. Interlayer tunneling spectroscopy is thus expected to display coherent peaks at half the driving frequency (alongside the interlayer-coherence-associated zero-bias peak). (c) In a Josephson-analogous geometry (weak links denoted by dashed black lines), the relative phase of the condensate between the driven part (driving represented by fuzzy orange oval) and the static part would result in an ac counterflow current $I\left(t\right)$ dominated by a $\omega_{d}/2$ harmonic.
@@ IV.2 Complex orders in a flat band model
The single-particle Hamiltonian of the two-dimensional (effectively spinless) system we consider here has valley $\tau_{i}$, pseudo-spin (or sub-lattice) $\sigma_{i}$, and orbital $\rho_{i}$ degrees of freedom (all Pauli matrices),
$H_{0}=-t\sum_{{\bf k}}c_{{\bf k}}^{\dagger}\left[\hat{h}_{x}\left({\bf k}\right)\rho_{x}+\hat{h}_{y}\left({\bf k}\right)\rho_{y}-\Delta_{\sigma}\sigma_{z}\right]c_{{\bf k}}.$ (53)
The matrices $\hat{h}_{x/y}\left({\bf k}\right)$ have internal structure in the $\sigma$-$\tau$ Hilbert space, and their exact form is detailed in Appendix F. Notably, $\left[\hat{h}_{x}\left({\bf k}\right)\right]^{2}+\left[\hat{h}_{y}\left({\bf k}\right)\right]^{2}=1$ for all $\bf k$, resulting in a spectrum of completely flat bands. The model is related to a tunable quantum metric model introduced in Refs. [110, 111]. The model is characterized by a parameter $\zeta$, corresponding to a quantum geometric length scale (in units of the unit cell) associated with the real-space spread of Wannier functions in the flat bands, referred to henceforth as the quantum metric length [112].
The Hamiltonian Eq. (53) has a total of eight identically flat bands: for $\Delta_{\sigma}=0$ (which we assume henceforth unless mentioned otherwise) one finds four valence bands sitting at energy t-t, and four conduction bands at +t+t. Our interest will lie in one set of these isolated bands, e.g., the valence bands. Within such set, reminiscent of the decomposition of flat bands in twisted graphene, one may define a good quantum number which further classifies the bands, $\sigma_{z}\tau_{z}$   [^4] In TBG, this matrix would correspond to the actual Chern number associated with each of these bands, yet in our model the Berry curvature vanishes everywhere – the model is topologically trivial, yet its quantum geometry is highly non-trivial for finite values of $\zeta$..
We examine interactions between the flat band electrons, projecting to the lowest lying four bands. The simplest density-density interaction term (that will dominate the interacting physics) $H_{\rm C}$, has the same form as in Eq. (43), except for modified density form factors,
$\Lambda_{{\bf kk^{\prime}}}={\cal F}_{{\bf kk^{\prime}}}+{\cal D}_{{\bf kk^{\prime}}}\sigma_{z}\tau_{z}.$ (54)
[…]
$\left[{\cal O}_{\rm g.s.},\sigma_{z}\tau_{z}\right]=0.$ (55)
Next, let us consider two sub-leading interaction terms, which favor a subset of the symmetry-breaking order parameters in Eq. (55),
@@
with $\rho^{\hat{M}}_{\bf q}=\sum_{\bf k}c_{{\bf k+q}}^{\dagger}\Lambda_{{\bf k+q,k}}\hat{M}c_{{\bf k}}$ and $u^{\rm K/T}_{\bf q}<0$.
The $H_{\rm K}$ interaction lowers the energy of a phase with finite expectation value of the operator $\sigma_{y}e^{i\theta\tau_{z}}\tau_{x}$, which corresponds to the celebrated Kramers intervalley coherent order (KIVC) in twisted bilayer graphene (TBG). This order has been broadly established as the ground state at half-filling in pristine TBG devices, both analytically and numerically [103, 104, 114]. This order parameter, however, is sensitive to the appearance of strain in realistic devices [115, 104, 116], often leading to the emergence of a rather different intervalley coherent phase, the so-called Kekulé spiral [117].
Still, even in nearly-pristine devices, with little to no strain, KIVC order is missing. Instead, a third intervalley coherent phase appears in experiments [118] – the time-reversal invariant intervalley coherent order (TIVC). In our model, this order is emulated by the order parameter $\sigma_{x}e^{i\theta\tau_{z}}\tau_{x}$, and is thus favored by the $H_{\rm T}$ interaction Hamiltonian. It has been suggested that electron-phonon interactions involving exchange of valley momentum favor the formation of the TIVC state over KIVC [82]. To capture the essence of the competition between these phases we consider the interactions in (56)–(57) short-ranged ($u^{\rm K/T}_{\bf q}$ are independent of the subscript $\bf q$, henceforth omitted), and define
$u\equiv-\left(u^{\rm K}+u^{\rm T}\right),$
$\Delta_{\rm TK}\equiv u^{{\rm K}}-u^{{\rm T}},$
$u_{\zeta}\equiv\left[1-J_{0}^{4}\left(2\zeta\right)\right].$
The TIVC phase is thus favored when $\Delta_{\rm TK}>0$. The parameter $u_{\zeta}$ is related to the energy difference between the lower energy KIVC and TIVC phases and the rest of the Eq. (55) manifold. Its dependence on quantum geometric length scale $\zeta$ stems from the form factors in Eq. (54), detailed in Appendix F. We henceforth assume that $u_{\zeta}>\Delta_{\rm TK}$, faithfully capturing the hierarchy of competing ordered ground states (see for example Fig. 6a).
Finally, alignment of the twisted graphene devices to the underlying hexagonal Boron Nitride (hBN) substrate breaks the inversion symmetry. This effect is approximately captured by in our model by a single-particle sublattice polarization term in Eq. (53), $\sim\Delta_{\sigma}\sigma_{z}$ [119, 120]. If this term is large enough, a sublattice-polarized phase, with an order parameter corresponding to $\sigma_{z}$, will be the energetically favored ground-state.